# QuiZSF: An efficient data-model interaction framework for zero-shot time-series forecasting

### v1

@@ 4.3 Hybrid and Hierarchical Time-series Retrieval
To support both high-accuracy and scalable search, we propose a hybrid and hierarchical TS retrieval (HHTR) strategy. It combines local domain-specific matching and global prototype comparison, leveraging the hierarchical index structure built in CRB. The strategy is shown in Figure 3 (a).
@@
Top-K Series Retrieval. Given a query time series $\mathbf{X}_{T}$, the objective is to retrieve its most relevant neighbors from the prototype database. This is a classical nearest-neighbor retrieval task, tailored to the time-series domain through hybrid matching strategies. If the domain of $\mathbf{X}_{T}$ is known and exists in CRB, we apply a combination of local and global prototype matching. The final Top-$K$ set is computed as,
$\text{Top}\;\;K=\rho\cdot\text{Top}\;\;K_{\text{local}}+(1-\rho)\cdot\text{Top}\;\;K_{\text{global}},$ (2)
@@
where $\rho\in[0,1]$ controls the balance between local and global contributions. If the domain is not identified or not present in CRB, retrieval is performed across all stored cluster prototypes,
$\text{Top}\;\;K=\text{Top}\;\;K_{\text{global}},$ (3)
This hybrid design enhances both retrieval accuracy and generalization, especially in settings with noisy or ambiguous domain information.
@@
$\text{Sim}(\mathbf{X}_{T},\mathbf{X}_{i})=\cos(\mathbf{X}_{T},\mathbf{X}_{i})+\frac{1}{\text{dist}(\mathbf{X}_{T},\mathbf{X}_{i})},$ (4)
@@
Figure 3: (a) HHTR: Domain-aware and global retrieval via a hierarchical index. (b) MSIL: Interaction and average patterns extracted from retrieved sequences are fused with the target via cross-attention.
4.4 Multi-grained Series Interaction Learner
The core challenge after retrieving relevant series lies in effectively integrating them with the target series to enhance forecasting. Simply concatenating or averaging the retrieved series can lead to suboptimal results due to heterogeneity across domains. To this end, we propose the Multi-grained Series Interaction Learner (MSIL), which is designed to extract robust and informative representations by modeling fine-grained interactions and global trends simultaneously.
MSIL is motivated by three considerations: (1) retrieved series may come from different domains with distinct dynamics; (2) fine-grained dependencies across series are often crucial for prediction; and (3) domain-specific context and global knowledge should be fused in a unified representation. As illustrated in Figure 3 (b), MSIL achieves this by computing two representative patterns, namely an interaction pattern and an average pattern, which are then fused with the target sequence through a cross-attention mechanism.
@@ 4.4 Multi-grained Series Interaction Learner
$\mathbb{S}_{\text{same}}=\{\mathbf{S}_{i}\mid\text{domain}(\mathbf{S}_{i})=\text{domain}(\mathbf{T})\},\quad\mathbb{S}_{\text{cross}}=\{\mathbf{S}_{j}\mid\text{domain}(\mathbf{S}_{j})\neq\text{domain}(\mathbf{T})\}.$ (5)
To ensure numerical consistency, we normalize both target and retrieved sequences using the scaler module associated with their respective base models,
$\mathbf{T}^{\text{norm}},\text{loc}_{T},\text{scale}_{T}$ $=\text{scaler}(\mathbf{T}),$ $\mathbf{S}_{i}^{\text{norm}},\text{loc}_{S_{i}},\text{scale}_{S_{i}}$ $=\text{scaler}(\mathbf{S}_{i}),\quad\forall i,$ (6)
@@
$\mathbf{T}^{\text{norm}}=\frac{\mathbf{T}-\text{loc}_{T}}{\text{scale}_{T}},\quad\mathbf{S}_{i}^{\text{norm}}=\frac{\mathbf{S}_{i}-\text{loc}_{S_{i}}}{\text{scale}_{S_{i}}}.$ (7)
Based on the normalized retrieved sequences, MSIL computes two complementary patterns: 1) Interaction Pattern ($\mathbf{P}_{\text{int}}$) captures fine-grained dependencies via element-wise product, followed by a non-linear projection. 2) Average Pattern ($\mathbf{P}_{\text{avg}}$) encodes global trends through mean pooling and transformation. These are defined as,
$\mathbf{P}_{\text{int}}$ $=\text{MLP}_{1}\left(\frac{\prod_{i=1}^{n}\mathbf{S}_{i}^{\text{norm}}}{\left\|\prod_{i=1}^{n}\mathbf{S}_{i}^{\text{norm}}\right\|}\right),$ $\quad\mathbf{P}_{\text{avg}}$ $=\text{MLP}_{2}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{S}_{i}^{\text{norm}}\right).$ (8)
To fuse these patterns with the target sequence, we use a multi-grained cross-attention mechanism, where the target serves as query and the patterns as key/value,
$\mathbf{Q}=\mathbf{W}_{q}\mathbf{T}^{\text{norm}},\quad\mathbf{K}=\mathbf{W}_{k}\mathbf{P}_{\text{avg}},\quad\mathbf{V}=\mathbf{W}_{v}\mathbf{P}_{\text{int}},$ (9)
$\mathbf{R}_{\text{fused}}=\text{Softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{T}}{\sqrt{d}}\right)\cdot\mathbf{V}.$ (10)
@@
4.5 Model Cooperation Coherer
In retrieval-augmented zero-shot forecasting, it is crucial to effectively connect retrieved knowledge with diverse TSPMs. These models fall into two categories: Non-LLM-based (numerical input) and LLM-based (textual input), each requiring tailored integration strategies. To fully leverage the representations produced by MSIL, we design a unified cooperation mechanism with two branches: one for numerical models and one for language models, both supporting feedback-driven optimization.
Numerical Coherer for Non-LLM based TSPMs. For numerical TS pre-trained models, we apply a residual connection to fuse the normalized target sequence $\mathbf{T}^{\text{norm}}$ with the MSIL-fused representation $\mathbf{R}_{\text{fused}}$, then feed it into the forecasting model,
$\widehat{\mathbf{T}}=\mathcal{F}_{\text{N}}(\mathbf{R}_{\text{fused}},\mathbf{T}^{\text{norm}}),$ (11)
where $\mathcal{F}_{\text{N}}$ is a residual module he2016deep that enhances expressiveness and mitigates gradient vanishing. The output $\widehat{\mathbf{T}}$ is the forecasted sequence.
MMD-based Optimization Strategy. To further align the predicted and ground truth distributions, we introduce a regularization term based on Maximum Mean Discrepancy (MMD) borgwardt2006integrating . Formally, given predicted sequences $\widehat{\mathbf{T}}\sim\mathcal{P}_{\widehat{T}}$ and ground truth sequences $\mathbf{T}_{\text{true}}\sim\mathcal{P}_{T}$, the MMD is defined as:
$\text{MMD}^{2}(\mathcal{P}_{\widehat{T}},\mathcal{P}_{T})=\left\|\mathbb{E}_{\widehat{\mathbf{T}}}[\phi(\widehat{\mathbf{T}})]-\mathbb{E}_{\mathbf{T}_{\text{true}}}[\phi(\mathbf{T}_{\text{true}})]\right\|_{\mathcal{H}}^{2},$ (12)
where $\phi(\cdot)$ is a feature map that embeds time series into a Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$, and $\|\cdot\|_{\mathcal{H}}$ denotes the RKHS norm. Using the kernel trick with $k(\mathbf{x},\mathbf{y})=\langle\phi(\mathbf{x}),\phi(\mathbf{y})\rangle$, MMD can be computed as:
$\text{MMD}^{2}=\mathbb{E}_{\widehat{\mathbf{T}},\widehat{\mathbf{T}}^{\prime}}[k(\widehat{\mathbf{T}},\widehat{\mathbf{T}}^{\prime})]+\mathbb{E}_{\mathbf{T}_{\text{true}},\mathbf{T}_{\text{true}}^{\prime}}[k(\mathbf{T}_{\text{true}},\mathbf{T}_{\text{true}}^{\prime})]-2\mathbb{E}_{\widehat{\mathbf{T}},\mathbf{T}_{\text{true}}}[k(\widehat{\mathbf{T}},\mathbf{T}_{\text{true}})].$ (13)
This formulation encourages the distribution of predicted sequences to match that of the true sequences. We integrate this into the training objective as a regularizer:
$\mathcal{L}=\mathcal{L}_{\text{MSE}}(\widehat{\mathbf{T}},\mathbf{T}_{\text{true}})+\lambda\cdot\text{MMD}^{2}(\widehat{\mathbf{T}},\mathbf{T}_{\text{true}}),$ (14)
where $\lambda$ is a balancing hyperparameter. This loss is backpropagated to update the learnable parameters, enhancing distributional alignment between prediction and truth.
5 Experiments
5.1 Experiments Setups
For evaluation datasets, we use five public datasets: ETTh1, ETTh2, ETTm1, ETTm2, and Weather, which are widely used in prior state-of-the-art works nie2022time ; jin2023time ; zhou2023one . Standard error metric is MSE. For CRB selection, our ChronoRAG Base comes in three versions: CRB-Small, CRB-Medium, and CRB-Large, with details in Appendix A.3. Given the balance between computing time and experimental results, we select CRB-Medium as the RAG Base. For model comparison, we evaluate against thirteen state-of-the-art forecasting methods, which can be classified into the following categories: (a) Non-LLM based TSPMs: TTM ekambaram2024tiny , Moirai woo2024unified , and TimesFM das2024decoder ; (b) LLM based TSPMs: TimeLLM jin2023time , LLMTime gruver2024large , and GPT4TS zhou2023one ; (c) Other architectures: iTransformer liu2023itransformer , Crossformer zhang2023crossformer , DLinear zeng2023transformers , TimesNet wu2022timesnet , PatchTST nie2022time , TiDE das2023long , and FEDformer zhou2022fedformer .
5.2 Implementation Details
In the time series domain, zero-shot forecasting refers to evaluating models on unseen datasets without direct supervision or fine-tuning. Given the two types of TSPMs, current ZSF setups are categorized accordingly: the multi-source generalization setup for Non-LLM based models, and the single-source transfer setup for LLM based models. To ensure comprehensive evaluation, we adopt both setups in our experiments, resulting in QuiZSFT and QuiZSFL, respectively.
Multi-source generalization zero-shot setup (for Non-LLM based TSPMs) This setup trains on diverse source datasets and enables zero-shot forecasting by directly applying the model to unseen target datasets. This approach is widely used in Non-LLM based TSPMs like TTM ekambaram2024tiny , TimesFM das2024decoder , and Moirai woo2024unified . We use this setup for QuiZSFT, with TTM-Base ekambaram2024tiny as the base model. The training datasets includes 38.7 million time points from multiple public benchmarks. Evaluation is conducted on the held-out ETT and Weather datasets, ensuring strict zero-shot conditions without data leakage. Detailed information can be found in Figure 9 of Appendix.
Single-source transfer zero-shot setup (for LLM based TSPMs) This setup trains a model on a single source dataset (e.g., ETTh1) and evaluates it on an unseen target dataset (e.g., ETTm2), highlighting the model’s ability to transfer knowledge across domains. Commonly used in LLM based TSPMs jin2023time ; zhou2023one , this setup involves fine-tuning or prompting LLMs on a specific dataset and then applying them to novel inputs from a different domain. We use this setup for QuiZSFL, with TimeLLM jin2023time as the base model and LLaMA-7B touvron2023llama as the backbone. Prompts are constructed with MSIL and domain-specific patterns, and evaluation is conducted across multiple train–test splits according to the TimeLLM ZSF experimental setup (Table 2).
5.3 Peformance Comparison
QuiZSFT. To provide a more comprehensive assessment, we compare QuiZSFT with several strong full-shot forecasting models. These comparisons fall under the category of long-sequence forecasting. As shown in Table 1, "zero-shot" refers to the prediction results of various base models without any pre-training on the test datasets, while "full-shot" denotes the performance of benchmark models that have been fully trained on each dataset. QuiZSFT, trained solely under the zero-shot setting, outperforms not only existing zero-shot baselines but also full-shot models, demonstrating strong generalization capabilities even without access to target-domain training data. QuiZSFT ranks Top1 in 75% of ZSF settings. However, we observe that QuiZSFT performs particularly well on relatively coarse-grained datasets but shows limited effectiveness on short-term, minute-scale forecasting tasks (e.g., ETTm1). This may be due to the difficulty in aligning retrieved information with fine-grained fluctuations in the target sequence. Leveraging coarse-grained knowledge to guide fine-grained prediction presents an interesting direction for future research. Combining complexity and efficiency comparison in Figure 4 (a), we consider QuiZSFT to be an excellent lightweight zero-shot forecasting framework, which benefits from the retrieval-enhanced representation and active feedback.
@@ 5.3 Peformance Comparison
Table 1: Long sequence forecasting results. The experimental setup follows TTM ekambaram2024tiny . Best results are in bold; second best are underlined. Full-shot results are obtained from the Moirai woo2024unified where the authors draw similar comparison.
@@
Table 2: Zero-shot forecasting results under the single-source transfer setup. Following the setting of TimeLLM jin2023time , results are averaged over prediction lengths {96, 192, 336, 720}. Best scores are in bold, second best are underlined.
@@
QuiZSFL. As shown in Table 2, QuiZSFL, equipped with retrieved augmented series, outperforms the majority of competing methods, achieving the best results in 7 out of 8 prediction settings. Additionally, our approach enhances the performance based on the pre-trained model. With the continuous advancement of pre-trained models, it is anticipated that QuiZSFL will bring about further performance improvements when adapting to new model frameworks.
5.4 Ablation study
Ablative variants. 1) QuiZSFL-w/o-RAG. We remove the auxiliary sequence retrieval and only utilize the LLM for ZSF to verify the motivation for RAG, which degenerates to LLMTime (gruver2024large, ). 2) QuiZSFL-w/o-MSIL. We remove the Multi-grained Series Interaction Learner, performing only an average calculation on the retrieved time series instead of feature extraction. 3) QuiZSFL-w/o-Coherer. We remove the structured prompt template and directly concatenate MSIL-extracted features with the target sequence, without adapting them into LLM-compatible textual inputs.
Main results. Results are in Table 3. The most significant performance drop occurs when removing the retrieval module (RAG), confirming its importance in providing external contextual signals for zero-shot forecasting. Compared to the full model, this variant sees a performance decline of 3.14%–5.34% (line 1 vs. line 4). Removing MSIL and using simple averaging instead also degrades performance. This is because the model cannot learn multi-level representations and capture relationships between sequences, resulting in performance degradation (line 2 vs. line 4). Additionally, without converting numerical time series into LLM-understandable tokens using the prompt template, the LLM’s performance is inferior to the full QuiZSFL, with a drop of about 2% (line 3 vs. line 4). These results demonstrate the effectiveness of our integrated approach.
Table 3: Ablation studies on zero-shot settings
@@ 5.4 Ablation study
5.5 Detailed model analysis

### v2

@@ 4.2.3. Hybrid and Hierarchical Time-series Retrieval
To support both high-accuracy and scalable search, we propose a hybrid and hierarchical time-series retrieval (HHTR) strategy. It combines local domain-specific matching and global prototype comparison, leveraging the hierarchical index structure built in CRB. The strategy is shown in Figure 3 (a).
Figure 3. (a) HHTR: Domain-aware and global retrieval via a hierarchical index. (b) MSIL: Interaction and average patterns extracted from retrieved sequences are fused with the target via cross-attention.
@@
Top-K Series Retrieval. Given a query time series $\mathbf{X}_{T}$, the objective is to retrieve its most relevant neighbors from the prototype database. This is a classical nearest-neighbor retrieval task, tailored to the time-series domain through hybrid matching strategies. When the domain of $\mathbf{X}_{T}$ is known and exists in CRB, we apply a combination of local and global prototype matching. The final Top-$K$ set is computed as:
(2) $\text{Top}\;\;K=\rho\cdot\text{Top}\;\;K_{\text{local}}+(1-\rho)\cdot\text{Top}\;\;K_{\text{global}},$
@@
where $\rho\in[0,1]$ controls the balance between local and global contributions. When the domain is not identified or not present in CRB, retrieval is performed across all stored cluster prototypes,
(3) $\text{Top}\;\;K=\text{Top}\;\;K_{\text{global}},$
Compared with existing methods, the proposed hybrid design exhibits superior performance in retrieval accuracy while maintaining higher computational efficiency.
@@
(4) $\text{Sim}(\mathbf{X}_{T},\mathbf{X}_{i})=\cos(\mathbf{X}_{T},\mathbf{X}_{i})+\frac{1}{\text{dist}(\mathbf{X}_{T},\mathbf{X}_{i})},$
@@
4.3. Multi-grained Series Interaction Learner
The core challenge after retrieving relevant series lies in effectively integrating them with the target series to enhance forecasting performance. Simply concatenating or averaging the retrieved series can lead to suboptimal results due to heterogeneity across domains. To this end, we propose the Multi-grained Series Interaction Learner (MSIL), which is designed to extract robust and informative representations by modeling fine-grained interactions and global trends simultaneously.
MSIL is motivated by three considerations: (1) retrieved series may come from different domains with distinct dynamics; (2) fine-grained dependencies across series are often crucial for forecasting; and (3) domain-specific context and global knowledge should be fused in a unified representation. As illustrated in Figure 3 (b), MSIL achieves this by computing two representative patterns, namely an interaction pattern and an average pattern, which are then fused with the target sequence through a cross-attention mechanism.
@@ 4.3. Multi-grained Series Interaction Learner
(5) $\mathbf{S}_{i}\in\begin{cases}\mathbb{S}_{\text{same}}&\text{if }\text{domain}(\mathbf{S}_{i})=\text{domain}(\mathbf{T}),\\ \mathbb{S}_{\text{cross}}&\text{if }\text{domain}(\mathbf{S}_{i})\neq\text{domain}(\mathbf{T})\end{cases}\quad\forall i\in n$
To ensure numerical consistency, we normalize both target and retrieved sequences using the scaler module associated with their respective base models:
(6) $\mathbf{T}^{\text{norm}},\text{loc}_{T},\text{scale}_{T}$ $=\text{scaler}(\mathbf{T}),$
(7) $\mathbf{S}_{i}^{\text{norm}},\text{loc}_{S_{i}},\text{scale}_{S_{i}}$ $=\text{scaler}(\mathbf{S}_{i}),\quad\forall i=1,2,\ldots,K$
@@
(8) $\mathbf{T}^{\text{norm}}=\frac{\mathbf{T}-\text{loc}_{T}}{\text{scale}_{T}},\quad\mathbf{S}_{i}^{\text{norm}}=\frac{\mathbf{S}_{i}-\text{loc}_{S_{i}}}{\text{scale}_{S_{i}}}$
Based on the normalized retrieved sequences, MSIL computes two complementary patterns: 1) Interaction Pattern ($\mathbf{P}_{\text{int}}$) captures fine-grained dependencies via element-wise product, followed by a non-linear projection. 2) Average Pattern ($\mathbf{P}_{\text{avg}}$) encodes global trends through mean pooling and transformation. These are defined as:
(9) $\mathbf{P}_{\text{int}}$ $=\text{MLP}_{1}\left(\frac{\prod_{i=1}^{n}\mathbf{S}_{i}^{\text{norm}}}{\left\|\prod_{i=1}^{n}\mathbf{S}_{i}^{\text{norm}}\right\|}\right),$ $\quad\mathbf{P}_{\text{avg}}$ $=\text{MLP}_{2}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{S}_{i}^{\text{norm}}\right)$
To fuse these patterns with the target sequence, we use a multi-grained cross-attention mechanism, where the target serves as query and the patterns as key/value:
(10) $\mathbf{Q}=\mathbf{W}_{q}\mathbf{T}^{\text{norm}},\quad\mathbf{K}=\mathbf{W}_{k}\mathbf{P}_{\text{avg}},\quad\mathbf{V}=\mathbf{W}_{v}\mathbf{P}_{\text{int}},$
(11) $\mathbf{R}_{\text{fused}}=\text{Softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{T}}{\sqrt{d}}\right)\cdot\mathbf{V}$
@@
Table 2. Long sequence forecasting results. The experimental setup follows TTM (Ekambaram et al., 2024). Best results are in bold; second best are underlined. Full-shot results are obtained from the Moirai (Woo et al., 2024) where the authors draw similar comparison.
@@
4.4. Model Cooperation Coherer
In retrieval-augmented zero-shot time-series forecasting, it is crucial to effectively connect retrieved knowledge with diverse TSPMs. These models fall into two categories: Non-LLM-based (numerical input) and LLM-based (textual input), each requiring tailored integration strategies. To fully leverage the representations produced by MSIL, we design a unified cooperation mechanism with two branches: one for numerical models and one for language models, both supporting feedback-driven optimization.
Numerical Coherer for Non-LLM based TSPMs. For numerical TS pre-trained models, we apply a residual connection to fuse the normalized target sequence $\mathbf{T}^{\text{norm}}$ with the MSIL-fused representation $\mathbf{R}_{\text{fused}}$, then feed it into the forecasting model:
(12) $\widehat{\mathbf{T}}=\mathcal{F}_{\text{N}}(\mathbf{R}_{\text{fused}},\mathbf{T}^{\text{norm}})$
where $\mathcal{F}_{\text{N}}$ is a residual module (He et al., 2016) that enhances expressiveness and mitigates gradient vanishing, and the output $\widehat{\mathbf{T}}$ denotes the forecasted sequence.
5. Experiments
5.1. Experiments Setups
For evaluation datasets, we use five public datasets: ETTh1, ETTh2, ETTm1, ETTm2, and Weather, which are widely used in prior state-of-the-art works (Nie et al., 2022; Jin et al., 2023; Zhou et al., 2023). Standard error metric is MSE. For CRB selection, our ChronoRAG Base comes in three versions: CRB-Small, CRB-Medium, and CRB-Large. Given the balance between computing time and experimental results, we select CRB-Medium as the RAG Base. For model comparison, we evaluate against thirteen state-of-the-art forecasting methods, which can be classified into the following categories: (a) Non-LLM based TSPMs: TTM (Ekambaram et al., 2024), Moirai (Woo et al., 2024), and TimesFM (Das et al., 2024); (b) LLM based TSPMs: TimeLLM (Jin et al., 2023), LLMTime (Gruver et al., 2024), and GPT4TS (Zhou et al., 2023); (c) Other architectures: iTransformer (Liu et al., 2023), Crossformer (Zhang and Yan, 2023), DLinear (Zeng et al., 2023), TimesNet (Wu et al., 2022), TiDE (Das et al., 2023), PatchTST (Nie et al., 2022), and FEDformer (Zhou et al., 2022).
5.2. Implementation Details
In the time series domain, zero-shot time-series forecasting refers to evaluating models on unseen datasets without direct supervision or fine-tuning. Given the two types of TSPMs, current ZSF setups are categorized accordingly: the multi-source generalization setup for Non-LLM based models, and the single-source transfer setup for LLM based models. To ensure comprehensive evaluation, we adopt both setups in our experiments, resulting in QuiZSFT and QuiZSFL, respectively.
Multi-source generalization zero-shot setup (for Non-LLM based TSPMs) This setup trains on diverse source datasets and enables zero-shot forecasting by directly applying the model to unseen target datasets. This approach is widely used in Non-LLM based TSPMs like TTM (Ekambaram et al., 2024), TimesFM (Das et al., 2024), and Moirai (Woo et al., 2024). We use this setup for QuiZSFT, with TTM-Base (Ekambaram et al., 2024) as the base model. For training, we use a comprehensive dataset covering 38.7 million time points, curated from multiple public benchmarks to ensure diversity. Evaluation is conducted on the held-out ETT and Weather datasets, ensuring strict zero-shot conditions without data leakage. Details of the training datasets can be found in Appendix C.
Table 3. Zero-shot forecasting results under the single-source transfer setup. Following the setting of TimeLLM (Jin et al., 2023), results are averaged over prediction lengths {96, 192, 336, 720}. Best scores are in bold, second best are underlined.
@@ 5.2. Implementation Details
Table 4. Ablation studies on zero-shot settings
@@
Single-source transfer setup (for LLM based TSPMs) This setup trains a model on a single source dataset (e.g., ETTh1) and evaluates it on an unseen target dataset (e.g., ETTm2), highlighting the model’s ability to transfer knowledge across domains. Commonly used in LLM based TSPMs (Jin et al., 2023; Zhou et al., 2023), this setup involves fine-tuning or prompting LLMs on a specific dataset and then applying them to novel inputs from a different domain. We use this setup for QuiZSFL, with TimeLLM (Jin et al., 2023) as the base model and LLaMA-7B (Touvron et al., 2023) as the backbone. The evaluation is carried out in multiple train-test splits according to the experimental setup of TimeLLM zero-shot forecasting (Table 3).
5.3. Peformance Comparison
QuiZSFT. Currently, there is limited work on zero-shot forecasting. To provide a more comprehensive assessment, we compare QuiZSFT not only with zero-shot methods but also with several strong full-shot forecasting models. These comparisons fall under the category of long-sequence forecasting. As shown in Table 2, "zero-shot" refers to the prediction results of various base models without any pre-training on the test datasets, while "full-shot" denotes the performance of benchmark models that have been fully trained on each dataset. QuiZSFT, trained solely under the zero-shot setting, outperforms not only existing zero-shot baselines but also full-shot models, demonstrating strong generalization capabilities even without access to target-domain training data. QuiZSFT ranks Top1 in 75% of ZSF settings. However, we observe that QuiZSFT performs particularly well on relatively coarse-grained datasets but shows limited effectiveness on short-term, minute-scale forecasting tasks (e.g., ETTm1). This may be due to the difficulty in aligning retrieved information with fine-grained fluctuations in the target sequence. Leveraging coarse-grained knowledge to guide fine-grained prediction presents an interesting direction for future research. Combining complexity and efficiency comparison in Figure 4 (a), we consider QuiZSFT to be an excellent lightweight zero-shot forecasting framework, which benefits from the retrieval-enhanced representation and active feedback.
QuiZSFL. As shown in Table 3, QuiZSFL, equipped with retrieved augmented series, outperforms the majority of competing methods, achieving the best results in 7 out of 8 prediction settings. Additionally, our approach enhances the performance based on the pre-trained model. With the continuous advancement of pre-trained models, QuiZSFL will bring about further performance improvements when adapting to new model frameworks.
Figure 4. (a) Size and time overview of QuiZTST vs. pre-trained TS benchmarks. Plot each model based on model size and the CPU inference time per batch. (b) Hyperparameter analysis on ETTm1 -> ETTm2.
Figure 5. Case studies on ETTh2 prediction.
5.4. Ablation study
Ablative variants. 1) QuiZSFL-w/o-RAG. We remove the auxiliary sequence retrieval and only utilize the LLM for ZSF to verify the motivation for RAG, which degenerates to LLMTime (Gruver et al., 2024). 2) QuiZSFL-w/o-MSIL. We remove the Multi-grained Series Interaction Learner, performing only an average calculation on the retrieved time series instead of feature extraction. 3) QuiZSFL-w/o-Coherer. We remove the structured prompt template and directly concatenate MSIL-extracted features with the target sequence, without adapting them into LLM-compatible textual inputs.
Main results. Results are presented in Table 4, with key findings summarized as follows: (i) The most significant performance drop occurs when removing the retrieval module (RAG), confirming its importance in providing external contextual signals for zero-shot forecasting. Compared to the full model, this variant sees a performance decline of 3.14%–5.34% (line 1 vs. line 4). (ii) Removing MSIL and using simple averaging instead also degrades performance. This is because the model cannot learn multi-level representations and capture relationships between sequences, resulting in performance degradation (line 2 vs. line 4). (iii) Additionally, without converting numerical time series into LLM-understandable tokens using the prompt template, the LLM’s performance is inferior to the full QuiZSFL, with a drop of about 2% (line 3 vs. line 4). These results demonstrate the effectiveness of our integrated approach.
5.5. Detailed model analysis